Copy Cats charges a $21 surcharge for a batch of colored copies plus $0.15 per page. Heritage charges a $25.50 surcharge for a batch of colored copies plus $0.12 per page. How many pages of color copies would make Heritage cheaper than Copy cats?
Answer by rfer(16322)

21+0.15x=25.50+0.12x
0.03x=4.5
x=4.5/0.03
x=150
more than 150 copies